the permiter of a rectangle is 54 cm and the base length is twice the height. What is the area
Answer by checkley79(3341)

P=2L+2H
L=2H
54=2*2H+2H
54=4H+2H
54=6H
H=54/6
H=9 ANS. FOR THE HEIGHT.
L=2*9=18 ANS. FOR THE LENGTH.
PROOF:
54=2*18=2*9
54=36+18
54=54
AREA=LW
18*9=162 CM^2 IS THE AREA.